# Geometric and topological rigidity of pinched submanifolds II

### v1

@@ Abstract
We continue the study of the geometry and topology of compact submanifolds of arbitrary codimension in space forms satisfying a certain pinching condition involving the length of the second fundamental form and the mean curvature. Our primary focus is on four-dimensional submanifolds, where, to our surprise, both the results obtained and the methods employed differ significantly and are notably more intricate compared to those in higher dimensions. This study heavily relies on concepts from four-dimensional geometry, the geometry of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing a crucial role. The results are sharp and extend previous results by several authors, without imposing any further assumption on either the mean curvature or the fundamental group of the submanifold.
@@ 1 Introduction
A fundamental problem in differential geometry is understanding the interplay between the geometry and topology of Riemannian manifolds. In the context of submanifold theory, it is particularly intriguing to investigate how the geometry or topology of submanifolds in space forms is influenced by pinching conditions involving either intrinsic or extrinsic curvature invariants.
For minimal submanifolds of spheres with a sufficiently pinched second fundamental form, an important result was first established by Simons in his seminal paper [22]. Subsequently, Chern, do Carmo, and Kobayashi [4] proved a celebrated rigidity theorem. Their work has since inspired numerous significant results in the study of pinching conditions. Notable examples include [1, 9, 12, 17, 21, 23, 28].
In our previous paper [24], we investigated the geometric and topological rigidity of submanifolds $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ satisfying the pinching condition
@@
pointwise. Here, $k$ is an integer with $21\leq k\leq n/$ 2, and the function $a$ is defined as
@@
Furthermore, $S$ denotes the squared length of the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$, which takes values in the normal bundle $N_{f}M$, while the mean curvature is defined as the length $H=\|\mathcal{H}\|$ of the mean curvature vector field given by $\mathcal{H}=(\mathrm{tr}\,\alpha_{f})/n$, where $\mathrm{tr}$ denotes the trace. The ambient space $\mathbb{Q}_{c}^{n+m}$ is the $(n+m)$-dimensional, complete, simply connected space form of constant curvature $c$. For simplicity, we assume that $c\in\{0,1\}$, unless stated otherwise. Thus $\mathbb{Q}_{c}^{n+m}$ is the Euclidean space $\mathbb{R}^{n+m}$ ($c=0$), or the unit sphere $\mathbb{S}^{n+m}$ ($c=1$).
A direct computation shows that the standard embedding of a torus
@@
into the unit sphere $\mathbb{S}^{n+1}$, where $\mathbb{S}^{k}(r)$ denotes the $k$-dimensional sphere of radius $r<1$, satisfies ($\ast$1) as an equality if $r\geq\sqrt{k/n}$, or $n=2k$. Otherwise, $S>a(n,k,H,1)$.
The pinching condition ($\ast$1) has been primarily studied for the lowest allowed $k$ (see for instance, [25, 27, 28, 30]). Shiohama and Xu [21] proved that compact submanifolds in space forms of nonnegative curvature are homeomorphic to a sphere, provided that ($\ast$1) holds as a strict inequality at any point for $k=1$. In [23], a homology vanishing result was established for submanifolds in space forms of nonnegative curvature that satisfy the strict inequality in ($\ast$1) pointwise for an integer $11\leq k\leq n-$ 1. The approach in that work relied on the nonexistence of stable currents, as shown by Lawson and Simons [10], under specific upper bounds of the second fundamental form.
In our recent paper [24], we investigated the geometric and topological rigidity of submanifolds of dimension $n\geq 5$ that satisfy the pinching condition ($\ast$1), without requiring the condition to be strict or imposing any assumptions on the mean curvature. We proved the following result, which shows that the pinching condition either enforces the vanishing of homology in a range of intermediate dimensions or uniquely determines the pinched submanifold up to congruence.
@@
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 5,c\geq 0$, be a substantial isometric immersion of a compact oriented Riemannian manifold. Assume that the inequality (italic-$\ast$1) is satisfied for an integer $22\leq k\leq n/$ 2 at any point. Then, either the homology groups of $M^{n}$ satisfy
@@
where $\beta_{k-1}(M)$ is the $(k-1)$-th Betti number, or equality holds in (italic-$\ast$1) at any point and one of the following assertions holds:
@@
This paper focuses on investigating the case of four-dimensional submanifolds, which was not addressed in Theorem 1. As will become clear, both the results obtained and the methods utilized differ substantially and are significantly more intricate than those used in the proof of Theorem 1 for higher dimensions. This study heavily relies on concepts from four-dimensional geometry, the geometry of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing an essential role. A key observation is that four-dimensional submanifolds satisfying the pinching condition ($\ast$1) possess nonnegative isotropic curvature, a concept first introduced by Micallef and Moore [14].
Throughout the paper, all submanifolds under consideration are assumed to be connected. We recall that a submanifold is called substantial if it is not contained in a proper totally geodesic submanifold of the ambient space. The main result of the paper can be stated as follows.
@@
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian four-manifold. Suppose the inequality (italic-$\ast$1) is satisfied for $k=2$ at every point. Then, one of the following assertions holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product ×$\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature, or
$(iii)$ Equality holds in (italic-$\ast$1) for $k=2$ at every point, and one of the following holds:
@@
    $M^{4}$ is isometric to the projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ of constant holomorphic curvature $4/3r^{2}$ with $r=1/\sqrt{c+H^{2}}$ and $f=j\circ g$, where $g$ is the standard embedding of $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ into $\mathbb{S}^{7}(r)$, and $j\colon\mathbb{S}^{7}(r)\to\mathbb{Q}_{c}^{8}$ is an umbilical inclusion.
@@
The standard isometric embedding of the torus $\mathbb{S}^{1}(r)\times\mathbb{S}^{3}(\sqrt{1-r^{2}})$ into $\mathbb{S}^{5}$ clearly satisfies ($\ast$1) for $k=2$, provided $r\geq 1/2$, and its fundamental group is infinite. In the final section, we present a method for constructing geometrically distinct isometric immersions of manifolds that are diffeomorphic to either $\mathbb{S}^{4}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{3}$, while also satisfying ($\ast$1) for $k=2$. These examples demonstrate that Theorem 2 is indeed sharp.
For submanifolds of arbitrary dimension that satisfy ($\ast$1) for $k=2$, we prove the following result.
Corollary 3.
@@
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian manifold. Assume that the inequality (italic-$\ast$1) is satisfied for $k=2$ at every point. Then, one of the following holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or
@@
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product ×$\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature, or
$(iii)$ If $n\geq 5$, the homology of $M^{n}$ is given by
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for}\;\,2\leq p\leq n-2\;\,\text{and}\;\,H_{p}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}\;\,\text{for}\;\,p=1,n-1,$
@@
or equality holds in (italic-$\ast$1) for $k=2$ at every point, and the submanifold is as described in parts $(i)-(iii)$ of Theorem 1, or as in part $(iii)$ of Theorem 2.
The results presented here extend the findings of previous studies [1, 2, 12, 28, 29, 30], without requiring additional assumptions on either the mean curvature or the fundamental group of the submanifold. Furthermore, the examples provided in the final section demonstrate the sharpness of Corollary 3.
@@
Submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) for $k=1$ were studied in [24]. Together with the above results, this provides a comprehensive understanding of submanifolds satisfying ($\ast$1) for the two smallest admissible values of the integer $k$.
@@ 2.1 Geometry of 4-dimensional manifolds
Let $(M,{\langle}\cdot,\cdot{\rangle})$ be an oriented Riemannian manifold of dimension $n=4$ with Levi-Civita connection $\nabla$ and curvature tensor $R$ given by
@@
[…]
@@
The Bochner-Weitzenböck operator is a self-adjoint operator. If $X\in\Lambda^{2}T_{x}M$, the dual 2-form $\omega$ is defined by $\omega(v,w)={\langle}{\langle}X,v\wedge w{\rangle}{\rangle}$, and we may consider $X$ as the skew-symmetric endomorphism of the tangent space at $x$ by ${\langle}X(u),v{\rangle}={\langle}{\langle}X,v\wedge w{\rangle}{\rangle}$.
[…]
@@
Then the Bochner-Weitzenböck operator acts on $\Lambda^{2}TM$ by $\omega(\mathcal{B}^{[2]}(X_{1}\wedge X_{2}))=\mathcal{B}^{[2]}(\omega)(X_{1},X_{2})$. Taking $\omega$ to be the dual form to $w_{1}\wedge w_{2}$ yields (2.1). Throughout the paper, we will mostly identify 2-forms with 2-vectors.
@@
of the eigenspaces of the Hodge star operator $\ast\colon\Omega^{2}(M)\to\Omega^{2}(M)$. The sections of $\Omega_{+}^{2}(M)$ are called self-dual 2-forms, whereas the ones of $\Omega_{-}^{2}(M)$ are called anti-self-dual 2-forms. Accordingly, we have the splitting
$\Lambda^{2}T_{x}M=\Lambda_{+}^{2}T_{x}M\oplus\Lambda_{-}^{2}T_{x}M$
at any point $x$, where $\Lambda_{\pm}^{2}T_{x}M$ are the eigenspaces of the Hodge star operator $\ast\colon\Lambda^{2}T_{x}M\to\Lambda^{2}T_{x}M$. Both spaces $\Lambda_{+}^{2}T_{x}M$ and $\Lambda_{-}^{2}T_{x}M$ are $\mathcal{B}^{[2]}$-invariant (see [19, Prop. 1]). Then we have accordingly the decomposition $\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}$.
Suppose now that $M$ is a compact oriented Riemannian four-manifold. The Hodge theorem guarantees that every de Rham cohomology class on $M$ has a unique harmonic representative. In particular, the space $\mathscr{H}^{2}(M)$ of harmonic 2-forms decomposes as
@@
where +$\mathscr{H}^{2}_{+}(M)$ and $\mathscr{H}^{2}_{-}(M)$ are the spaces of self-dual and anti-self-dual harmonic 2-forms, respectively. The dimensions of these subspaces, denoted by $\beta_{\pm}(M)=\dim\mathscr{H}^{2}_{\pm}(M)$, are oriented homotopy invariants of $M$. Their difference $\sigma=\beta_{+}(M)-\beta_{-}(M)$ is the signature of $M$, while their sum equals the second Betti number $\beta_{2}(M)$ of the manifold $M$.

### v2

@@ Abstract
We continue the study of the geometry and topology of compact submanifolds of arbitrary codimension in space forms that satisfy a pinching condition involving the length of the second fundamental form and the mean curvature. Our primary focus is on four-dimensional submanifolds, where both the results obtained and the methods employed differ substantially and are considerably more intricate than in higher dimensions. This study relies critically on concepts from four-dimensional geometry, the theory of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing an essential role. The results are sharp and extend previous work by several authors, without imposing additional assumptions on either the mean curvature or the fundamental group of the submanifold.
@@ 1 Introduction
A fundamental problem in differential geometry is to understand the interplay between the geometry and topology of Riemannian manifolds. In the context of submanifold theory, it is particularly interesting to investigate how the geometry and topology of submanifolds in space forms are influenced by pinching conditions involving either intrinsic or extrinsic curvature invariants.
For minimal submanifolds of spheres with a sufficiently pinched second fundamental form, a fundamental result was first established by Simons in his seminal work [26]. Later, Chern, do Carmo, and Kobayashi [4] proved a celebrated rigidity theorem. Their contributions have since inspired a wealth of important developments in the study of pinching conditions, including, for example, [1, 11, 14, 20, 25, 27, 32].
In our previous paper [28], we studied the geometric and topological rigidity of submanifolds $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ satisfying the pointwise pinching condition
@@
where $S$ denotes the squared length of the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$, taking values in the normal bundle $N_{f}M$. The mean curvature is defined as $H=\|\mathcal{H}\|$, the norm of the mean curvature vector field
$\mathcal{H}=\frac{1}{n}\,\mathrm{tr}\,\alpha_{f},$
where $\mathrm{tr}$ denotes the trace. Here $k$ is an integer with $1\leq k\leq n/2$, and the function $a$ is defined as
@@
The ambient space $\mathbb{Q}_{c}^{n+m}$ is the $(n+m)$-dimensional, complete, simply connected space form of constant curvature $c$. For simplicity, we restrict to the cases $c\in\{0,1\}$, unless stated otherwise. Accordingly, $\mathbb{Q}_{c}^{n+m}$ is either the Euclidean space $\mathbb{R}^{n+m}$ when $c=0$, or the unit sphere $\mathbb{S}^{n+m}$ when $c=1$.
A straightforward computation shows that the standard embedding of a torus
@@
into the unit sphere $\mathbb{S}^{n+1}$, where $\mathbb{S}^{k}(r)$ denotes the $k$-dimensional sphere of radius $r<1$, satisfies ($\ast$1) with equality if $r\geq\sqrt{k/n}$, or $n=2k$. In all other cases, one has $S>a(n,k,H,1)$.
The pinching condition ($\ast$1) has been studied primarily in the case of the smallest admissible value of $k$ (see, e.g., [29, 31, 32, 34]). Shiohama and Xu [25] proved that compact submanifolds in space forms of nonnegative curvature are homeomorphic to a sphere, provided ($\ast$1) holds as a strict inequality at every point when $k=1$. In [27], a homology vanishing result was established for submanifolds in space forms of nonnegative curvature that satisfy the strict pointwise inequality in ($\ast$1) for some integer $1\leq k\leq n-1$. The argument there relied on the nonexistence of stable currents, as established by Lawson and Simons [12], under suitable upper bounds for the second fundamental form.
In our recent work [28], we studied the geometric and topological rigidity of submanifolds of dimension $n\geq 5$ satisfying the pinching condition ($\ast$1), without assuming strictness or imposing any restrictions on the mean curvature. Among other results, we established the following theorem, showing that the pinching condition either forces the vanishing of homology in certain intermediate dimensions or uniquely determines the submanifold up to congruence. Recall that a submanifold is said to be substantial if it is not contained in any proper totally geodesic submanifold of the ambient space.
@@
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 5,c\geq 0$, be a substantial isometric immersion of a compact oriented Riemannian manifold. Assume that inequality ($\ast$1) is satisfied for some integer integer $2\leq k\leq n/2$ at every point. Then, either the homology groups of $M^{n}$ satisfy
@@
where $\beta_{k-1}(M)$ is the (k1)(k-1)-th Betti number, or equality holds in ($\ast$1) at every point and one of the following assertions holds:
@@
The main focus of this paper is the study of four-dimensional submanifolds satisfying condition ($\ast$1), which are not covered by Theorem 1. The results obtained here, as well as the methods employed, differ substantially from those in the proof of Theorem 1 for higher dimensions and are considerably more intricate. Our approach relies essentially on tools from four-dimensional geometry, the theory of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique. A key observation is that four-dimensional submanifolds satisfying the pinching condition ($\ast$1) necessarily have nonnegative isotropic curvature, a concept introduced and studied by Micallef and Moore [17].
Throughout the paper, all submanifolds under consideration are assumed to be connected. The main result of the paper can be stated as follows.
@@
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian four-manifold. Suppose inequality ($\ast$1) is satisfied for $k=2$ at every point. Then, one of the following assertions holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$.
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product $\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature.
$(iii)$ Equality holds in ($\ast$1) for $k=2$ at every point, and one of the following holds:
@@
    $M^{4}$ is isometric to the projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ of constant holomorphic curvature $4/3r^{2}$ with $r=1/\sqrt{c+H^{2}}$ and $f=j\circ g$, where $g$ is the standard minimal embedding of $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ into $\mathbb{S}^{7}(r)$, and $j\colon\mathbb{S}^{7}(r)\to\mathbb{Q}_{c}^{8}$ is an umbilical inclusion.
@@
In the final section of the paper, we present a method for constructing geometrically distinct isometric immersions of manifolds diffeomorphic either to $\mathbb{S}^{4}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{3}$, which also satisfy ($\ast$1) for $k=2$. These examples demonstrate that Theorem 2 is sharp. In particular, the standard isometric embedding of the torus $\mathbb{S}^{1}(r)\times\mathbb{S}^{3}(\sqrt{1-r^{2}})$ into $\mathbb{S}^{5}$ satisfies ($\ast$1) for $k=2$ whenever $r\geq 1/2$, and its fundamental group is infinite. This example shows that the class of submanifolds appearing in part $(ii)$ of Theorem 2 is nonempty.
For submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) with $k=2$, we establish the following result.
Theorem 3.
@@
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, with $n\geq 4$ and $c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian manifold. Assume that inequality ($\ast$1) is satisfied at every point for $k=2$. Then, one of the following holds:
$(i)$ The manifold $M^{n}$ is diffeomorphic to $\mathbb{S}^{4}$.
@@
$(ii)$ The universal cover of $M^{n}$ is isometric to a Riemannian product $\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature.
$(iii)$ If $n\geq 5$, the homology of $M^{n}$ satisfies
$H_{p}(M^{n};\mathbb{Z})=0\;\;\text{for all}\;\;2\leq p\leq n-2$
and
$H_{p}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}\;\;\text{for}\;\;p=1,n-1.$
@@
$(iv)$ Equality holds in condition ($\ast$1) for $k=2$ at every point, and the submanifold is as described in parts (i)(i)-$(iii)$ of Theorem 1, or as in part $(iii)$ of Theorem 2.
The results presented here extend those of previous studies [1, 2, 4, 11, 14, 26, 32, 33, 34] without imposing additional assumptions on either the mean curvature or the fundamental group of the submanifold. In the final section, we further construct geometrically distinct isometric immersions of manifolds diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, with $n\geq 4$, which also satisfy ($\ast$1) for $c=0$ and $k=2$. These examples illustrate the sharpness of Theorem 3. Finally, note that the homology described in part $(iii)$ of Theorem 3 coincides with that of the sphere $\mathbb{S}^{n}$ with $g=\beta_{1}(M)$ handles attached, or equivalently, with the homology of the connected sum $_{g}(\mathbb{S}^{1}\times\mathbb{S}^{n-1})$ consisting of $g$ summands.
@@
The geometric and topological rigidity of submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) for $k=1$ was studied in [28]. Combined with the results presented here, this provides a comprehensive understanding of compact submanifolds satisfying ($\ast$1) for the two smallest admissible values of the integer $k$.
It is worth noting that in the recent paper [15], the authors extended the pinching result of Simons [26] and that of Chern, do Carmo, and Kobayashi [4] to complete (not necessarily compact) minimal submanifolds in spheres. It would be of interest to extend the results of the present paper, as well as those in [28], to complete submanifolds in space forms, without imposing any restrictions on the mean curvature. Such a generalization could provide deeper insights into the geometry and topology of submanifolds satisfying the pinching condition ($\ast$1), beyond the minimal or compact setting.
@@ 2.1 Geometry of 4-dimensional manifolds
Let $(M,{\langle}\cdot,\cdot{\rangle})$ be an oriented Riemannian manifold of dimension $n=4$ with Levi-Civitá connection $\nabla$ and curvature tensor $R$ given by
@@
and then extend it linearly to all of $\Lambda^{2}T_{x}M$. Here $\langle$$\langle$$\cdot$$\cdot{\rangle}{\rangle}$ stands for the inner product of $\Lambda^{2}T_{x}M$ defined by
@@
The Bochner-Weitzenböck operator is a self-adjoint operator. If $Z\in\Lambda^{2}T_{x}M$, the dual 2-form $\omega$ is defined by
$\omega(v,w)={\langle}{\langle}Z,v\wedge w{\rangle}{\rangle},$
and we may regard $Z$ as the skew-symmetric endomorphism of the tangent space at $x$ via
${\langle}Z(v),w{\rangle}={\langle}{\langle}Z,v\wedge w{\rangle}{\rangle}.$
Clearly $\mathcal{B}^{[2]}$ can also be viewed as an endomorphism of the bundle $\Omega^{2}(M)$ of 2-forms of the manifold via the inner product $\langle$$\langle$$\cdot$$\cdot{\rangle}{\rangle}$. If $\omega$ is a 2-form, then $\mathcal{B}^{[2]}(\omega)$ is given by
@@
Then the Bochner-Weitzenböck operator acts on $\Lambda^{2}TM$ by
$\omega(\mathcal{B}^{[2]}(X_{1}\wedge X_{2}))=\mathcal{B}^{[2]}(\omega)(X_{1},X_{2}).$
Taking $\omega$ to be the dual form to $w_{1}\wedge w_{2}$ yields (2.1). Throughout the paper, we will generally identify 2-forms with 2-vectors.
@@
of the eigenspaces of the Hodge star operator
$\ast\colon\Omega^{2}(M)\to\Omega^{2}(M)$
The sections of $\Omega_{+}^{2}(M)$ are called self-dual 2-forms, whereas the ones of $\Omega_{-}^{2}(M)$ are called anti-self-dual 2-forms. Accordingly, at any point $x\in M$, we have the splitting
$\Lambda^{2}T_{x}M=\Lambda_{+}^{2}T_{x}M\oplus\Lambda_{-}^{2}T_{x}M,$
where $\Lambda_{\pm}^{2}T_{x}M$ are the eigenspaces of the Hodge star operator
$\ast\colon\Lambda^{2}T_{x}M\to\Lambda^{2}T_{x}M.$
Both spaces $\Lambda_{+}^{2}T_{x}M$ and $\Lambda_{-}^{2}T_{x}M$ are $\mathcal{B}^{[2]}$-invariant (see Proposition 1 in [23]). Hence, we obtain the corresponding decomposition
$\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}.$
Suppose now that $M$ is a compact, oriented Riemannian four-manifold. The Hodge theorem guarantees that every de Rham cohomology class on $M$ has a unique harmonic representative. In particular, the space $\mathscr{H}^{2}(M)$ of harmonic 2-forms decomposes as
@@
where $\mathscr{H}^{2}_{+}(M)$ and $\mathscr{H}^{2}_{-}(M)$ are the spaces of self-dual and anti-self-dual harmonic 2-forms, respectively. The dimensions of these subspaces, denoted by $\beta_{\pm}(M)=\dim\mathscr{H}^{2}_{\pm}(M)$, are oriented homotopy invariants of $M$. Their difference
$\sigma=\beta_{+}(M)-\beta_{-}(M)$
is the signature of $M$, while their sum equals the second Betti number $\beta_{2}(M)$ of the manifold.